What is the pH of a 3.3 x 10^-7 M NaOH solution? [_____________] = pH

Please write out all steps
Answer #1

NaOH dissociates in aqueous solution fully, as-

NaOH Na+ + OH​​​​​- .

So, 3.3×10-7 M NaOH dissociates to give 3.3×10-7 M OH​​​​​- .

So, [OH​​​​​-] = 3.3×10-7

In aqua solution solution : [H+] × [ OH​​​​​- ] = 1×10-14

Hence, [H+] = 1×10-14/ [OH​​​​​-]

So, [H+] = (1×10-14)/(3.3×10-7)

Or, [H+] = 3.03 × 10-8 .

We know, pH = -log[H+]

So, pH = - log (3.03×10-8)

Hence, pH = 7.52.